Given the following data for a stock: beta = 1.4; risk-free rate = 3%; market rate of return = 13%; and expected rate of return on the stock = 15%. Then the stock is:

below the Security Market Line

on the Security Market Line

above the Security Market Line

it cannot be determined

  1. You own a portfolio of two stocks, O and Q. Stock O is valued at $4,000 and has an expected return of 10.5 percent. Stock Q has an expected return of 18.7 percent. What is the expected return on the portfolio if the portfolio total value is $5,000?

    11.73 percent

    10.92 percent

    12.14 percent

    13.03 percent

  1. Cornel Co. has a beat of 1.4. If the risk-free rate is 3% and the market return is 12%, what is its cost of equity assume CAPM?

    18.0%

    16.5%

    15.6%

    17.6%
Answer #1

Question 1

using Capital Asset Pricing Model

Required Return = Rf + b ( Rm – Rf )

Where,

Rf – Risk free return = 3%

b – Beta = 1.4

Rm – Expected return on market portfolio = 13%

Required Return = 3+1.4*(13-3)

= 3+14

= 17.00%

expected rate of return on the stock = 15%

expected rate of return on the stock = 15% < Required Return =17%, stocks are below the Security Market Line

Note: As per HomeworkLib guidlines we are mandated to answer only the 1st one when multiple questions are posted.